# Using Knowledge Graphs to harvest datasets for efficient CLIP model training

### v1

@@ A Detailed results
B-16 $\star$ EntityNet (only living organisms) 11.3 10.2 51.4 39.2 46.1 62.1 25.9 35.1 26.8
B-32 EntityNet Everything 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 EntityNet (no living organisms) 9.0 22.5 39.2 28.0 33.6 37.7 29.1 32.0 7.5
B-32 EntityNet (only living organisms) 4.1 10.2 36.0 21.0 31.5 39.6 8.1 17.6 8.1
B-32 EntityNet (no attribute queries) 8.7 21.8 54.8 33.8 47.9 49.2 24.4 36.9 10.5
B-32 EntityNet 50% alt text 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 EntityNet 100% alt text 13.1 32.7 59.1 38.1 51.1 55.6 30.1 40.8 13.1
B-32 EntityNet 0% alt text 13.1 32.7 55.7 35.5 48.0 53.0 27.1 36.8 12.4
B-32 EntityNet Full size 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 EntityNet 1/2 size 6.6 16.4 54.1 33.6 47.2 51.0 24.4 36.5 8.9
B-32 EntityNet 1/4 size 3.3 8.2 45.2 25.7 39.1 39.9 18.1 25.6 5.9
B-32 EntityNet 1/8 size 1.6 4.1 33.3 17.7 28.5 28.7 12.0 15.3 4.2
B-32 EntityNet 1/16 size 0.8 2.0 19.9 10.0 17.2 17.6 6.8 5.7 2.8
B-32 EntityNet batch size 2,048 13.1 32.7 59.7 40.4 52.5 59.4 31.3 44.5 14.4
B-32 EntityNet batch size 4,096 13.1 32.7 60.9 41.0 53.0 60.0 32.1 44.9 14.9
B-32 EntityNet batch size 8,192 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 EntityNet batch size 16,384 13.2 32.7 60.5 39.7 53.2 58.0 30.6 42.9 13.8
B-32 EntityNet batch size 32,768 13.3 32.7 58.6 37.2 51.3 54.6 27.9 40.6 11.6
B Qualitative examples of our dataset
We show randomly sampled images and corresponding textual information of our dataset in Figures 4, 5, 6, and 7. We show an example of our text label sampling strategy in Table 10.
Refer to caption Entity + attribute running polecat A wonderful polecat in its woodland surroundings Polecats Unveiled: Sleek Predators in the Countryside (Mustela Putorius) - Glenlivet Wildlife Polecats Unveiled: Sleek Predators in the Countryside (Mustela Putorius) Black Polecat Photos and Premium High Res Pictures - Getty Images Do Cats Eat Ferrets – What You Should Know! – FAQcats.com Do Cats Eat Ferrets – What You Should Know!
Refer to caption Entity + attribute summer Canada goose Canada Geese Goose Branta - Free photo on Pixabay - Pixabay Canada Geese Goose Branta · Free photo on Pixabay Facts about geese Canada Geese Goose Branta Free Photo On Pixabay Pixabay, 45% OFF
Refer to caption Entity + attribute wild tortoise Greek Tortoise Testudo Graeca Hiding Shell Stock Photo 1425661328 | Shutterstock Elongated Tortoise Indotestudo Elongata Yellow Tortoise Stock Photo 1463951543 | Shutterstock
Refer to caption Entity Orbea decaisneana Orbea decaisneana subs. hesperidum f. cristata
Refer to caption Entity + attribute old walrus What Is A Walrus? What is a Walrus - Walrus Habitat and Behavior - Wild Focus Expeditions Portrait of an old bull walrus resting on his teeth, tooth walker
Figure 4: Randomly sampled images from the EntityNet living subset together with query type, search query, and alt texts (separated by new lines).
Refer to caption Entity Junín red squirrel Curious Eurasian Red Squirrel, Sciurus Vulgaris, Running and Jumping …
Refer to caption Entity chile pine Araucaria araucana - Wikipedia Araucaria araucana - Wikipedia, la enciclopedia libre Araucaria araucana - Wikipedia | Trees to plant, Denver botanic gardens … a tall tree in the middle of a forest Araucaria araucana - Wikipedia | Denver botanic gardens, Outdoor plants … Monkey Puzzle Plant main Monkey Puzzle Plant Care & Growing Basics: Water, Light, Soil, Propagation etc.
Refer to caption Entity Cedronella canariensis Cedronella canariensis
Refer to caption Entity Barbuda Warbler In the face of elite tourism projects, the Barbuda Warbler isn’t the only one that might lose its home Barbuda Warbler - Setophaga subita - Birds of the World / - Barbuda Warbler
Refer to caption Natural type + attribute long neck animal The Strange Elegance of the Giraffe-Necked Antelope | The Ark In Space
Figure 5: Randomly sampled images from the EntityNet living subset together with query type, search query, and alt texts (separated by new lines).
Refer to caption Natural type + attribute carpaccio with lemon Tuna Carpaccio with Fennel and Lemon Recipe - Great British Chefs Tuna carpaccio with fennel and lemon
Refer to caption Natural type + attribute index finger nail A Macro Of Index Finger Nail Stock Footage SBV-301021776 - Storyblocks
Refer to caption Entity akaogiite Akaogiite Cut Out Stock Images
Refer to caption Entity public bookcase manufacta est: public bookcase
Refer to caption Entity BK 117 Eurocopter-Kawasaki BK-117B-2 - DRF - Deutsche Rettungsflugwacht | Aviation Photo #1053819 | Airliners.net Eurocopter-Kawasaki BK-117B-2 - DRF - Deutsche Rettungsflugwacht
Figure 6: Randomly sampled images from the EntityNet non-living subset together with query type, search query, and alt texts (separated by new lines).
Refer to caption Entity cowboy boots Ariat Heritage Rough Stock Cowboy Boots - Square Toe - Country Outfitter
Refer to caption Entity + attribute nacre jewelry Genuine Nacre Necklace Pearl Jewelry Mother Of Pearls Beads | Etsy
Refer to caption Entity + attribute small artificial pond Small backyard pond decoration. Artificial pond in garden. Pool aquatic plants. Pond border decoration. photo Small backyard pond decoration. Artificial pond in garden. Pool aquatic plants. Pond border decoration. Small backyard pond decoration. Artificial pond in garden. Pool aquatic plants. Pond border decoration. 9562060 Stock Photo at Vecteezy Small backyard pond decoration. Artificial pond in garden. Pool aquatic plants. Pond border decoration. Free Photo
Refer to caption Entity + attribute berimbau at rest Premium AI Image | Best Berimbau With Handle Isolated On White Background
Refer to caption Entity + attribute assembled motherboard Red Team rocking: Build the ultimate AMD gaming PC | PCWorld amd fx cpu
Figure 7: Randomly sampled images from the EntityNet non-living subset together with query type, search query, and alt texts (separated by new lines).
@@ B Qualitative examples of our dataset
C Hyperparameter settings
We report hyperparameters for our experiments in Table 11. Similar to Li et al. Li et al. [17], we reduce the context size of the text encoder down from 77 to 32 to reduce VRAM and training time requirements. For a fair comparison with other CLIP models, we report all training cost and training duration as if the training was run at a context length of 77.
D Image Search APIs
Google The Google Image Search API is available via the Google Cloud Platform, and requires an existing programmable search engine to function. It returns up to 10 images per request and page with a limit of 10 pages, i.e., 100 images per query. It costs 5\$ per 1,000 API calls, resulting in costs of about 500\$ to download 1M images. We found the search results from the Google API to be quite different, and arguably worse, from the ones returned when using the regular Google image search. For all our API requests we set the parameter imgColorType to color, imgType to photo, lr to lang_en, and excludeTerms to drawing clipart illustration cartoon vector painting. This way we get mostly real-world images in our search results. We additionally add all aliases and the natural type of the sought entity to the orTerms parameter for entity and entity-attribute queries. Because the Google API returns only up to 10 images per request and page, we search for the following number of pages: 2 pages each for entity queries, 4 pages each for entity-attribute queries, and 10 pages each for all natural-type-attribute queries. We started our search with queries from the living subset on both the Google and Bing APIs. We found the quality and value-for-money ratio of the Bing API to be better, and therefore switched to only using Bing for all other queries.
E Querying entities with SPARQL
The SPARQL query over Wikidata used to harvest all entities under a super-entity is displayed in Figure 8. It returns a list of entities from a specified target domain as defined by one or more super-entities. The super-entities can be determined manually by searching for appropriate entities on the Wikidata website. For example, if we want to build a dataset about vehicles, we can set the super-entity to vehicle (Q42889), as done in Table 12. We list the super-entities we considered for our dataset and the relevant statistics in Table 13 for included and Table 14 for excluded super-entities.
Table 11: Hyperparameters used for training and finetuning.
Dataset Model Batch Size Max LR
Weight
Decay
Epochs
Warmup
eps Beta 1 Beta 2
CC12M ViT-B/32 8k 5e-4 0.2 18 2 1e-8 0.9 0.98
CC12M ViT-B/16 8k 5e-4 0.2 18 2 1e-8 0.9 0.98
Ours ViT-B/32 8k 5e-4 0.2 18 2 1e-8 0.9 0.98
Ours ViT-B/16 8k 5e-4 0.2 18 2 1e-8 0.9 0.98
Ours, finetuning ViT-B/32 32k 5e-5 0.2 18 2 1e-8 0.9 0.98
Ours, finetuning ViT-B/16 32k 5e-5 0.2 18 2 1e-8 0.9 0.98
PREFIX wdt: <http://www.wikidata.org/prop/direct/>PREFIX wd: <http://www.wikidata.org/entity/>PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>PREFIX schema: <http://schema.org/>PREFIX wikibase: <http://wikiba.se/ontology#>PREFIX skos: <http://www.w3.org/2004/02/skos/core#>SELECT DISTINCT  ?ent  ?label  ?desc  ?links  (GROUP_CONCAT(DISTINCT ?alias; SEPARATOR=";;;") AS ?aliases)WHERE {  VALUES ?typ { wd:Q42889 }  ?ent wdt:P279* ?typ .  ?ent rdfs:label ?label .  FILTER(LANG(?label) = "en")  ?ent ^schema:about/wikibase:sitelinks ?links .  FILTER(?links >= 5)  OPTIONAL {    ?ent schema:description ?desc .    FILTER(LANG(?desc) = "en")  }  OPTIONAL {    ?ent skos:altLabel ?alias .    FILTER(LANG(?alias) = "en")  }}GROUP BY ?ent ?label ?desc ?linksORDER BY DESC(?links)
@@ E Querying entities with SPARQL
Table 12: Vehicle entities and accompanying information as extracted from the Wikidata knowledge graph. Showing the first 5 and last 5 out of 17,015 entities. Note that we only collect entities with sitelinks $\geq 5$. The corresponding SPARQL query is shown in Figure 8.
@@
Q1420 motor car motorized road vehicle designed to carry one to eight people rather than primarily goods 237 auto / motor vehicle / motor cars / motorcar / cars / car / automobiles / automobile / autocar
Q11442 bicycle pedal-driven two-wheel vehicle 203 bike / Bicycles / cycle / pushbike / pedal cycle / pedal bike
Q197 airplane powered fixed-wing aircraft 196 airplane, aeroplane, plane / powered fixed-wing aircraft / planes / plane / aeroplane / fixed-wing powered aircraft / fixed-wing airplane / aeroplanes / fixed-wing aeroplane / airplanes
Q870 train form of rail transport consisting of a series of connected vehicles 193 rail-train / trains / railway train / railtrain / rail train / railroad train
Q11446 ship large buoyant watercraft 178 marine vessel / vessel / water vessel / ships
@@
Q7077241 Odakyu 20000 series RSE Japanese electric multiple unit trainset 5 RSE / Romancecar RSE / Resort Super Express / Odakyu Romancecar RSE / 20000 series
Q812263 Bavarian Pt 2/3 class of 97 German 2-4-0T locomotives 5 ÖBB 770 / DR Class 70.0 / DRG Class 70.0
@@
Q812260 Bavarian PtL 2/2 class of 6+29+13 German 0-4-0T locomotives 5 DB Class 98.3 / DRG Class 98.3 / ÖBB 688
@@
product Anything that can be offered to a market banh mi, navigation system, PlayStation 2 63,676 144,715
substance Any composed matter whose origin is either biological, chemical, or mineral solid lubricant, Chinese tea, eye cups 34,259 111,383
physical tool Physical item that can be used to achieve a goal Patient lift, police transport, instant camera 32,727 71,227
animal Kingdom of multicellular eukaryotic organisms saw-scaled viper, Sporathraupis cyanocephala, Rufous mouse-eared bat 28,000 76,408
plant Living thing in the kingdom of photosynthetic eukaryotes Whitebark Pine, Eucalyptus coccifera, wig knapweed 28,000 55,925
material Substance that can occur in different amounts, all with some similar [mixture of some] characteristics, and with which objects can be made dietary proteins, stone slab tomb, safflower oil 18,021 40,822
vehicle Mobile machine used for transport, whether it has an engine or not, including wheeled and tracked vehicles, air-, water-, and space-craft shipwrecks (objects), Evergreen A-class container ship, VTOL aircraft 17,015 37,849
geographical feature Components of planets that can be geographically located hydrothermal Vents, grooves, street lamp 8,683 19,030
food Any substance consumed to provide nutritional support for the body; form of energy stored in chemical coffee milk, tikka, Friesian Clove 8,464 15,332
architectural structure Human-designed and -made structure rock temples, summerhouse, house of worship 4,507 10,354
anatomical structure Entity with a single connected inherent 3d shape that’s created by coordinated expression of the organism’s own dna bronchi, maxillary wisdom tooth, turtle shell 4,394 9,999
facility Place, equipment, or service to support a specific function public toilet, automobile servicing shop, industrial park 2,767 6,740
physical activity Human physical activity consisting of voluntary bodily movement by skeletal muscles American rules football, archery, water-skiing 2,228 4,422
clothing Covering worn on the body blucher shoe, G-suit, one-piece swimsuit 1,929 4,313
building Structure, typically with a roof and walls, standing more or less permanently in one place shoestore, family restaurant, factory outlet 1,655 3,964
musical instrument Device created or adapted to make musical sounds electroencephalophone, Chinese flutes, oboe 1,450 3,493
organ Collection of tissues with similar functions nasal bone, cranial nerves, ulnar collateral ligament of elbow 1,155 2,450
furniture Movable objects used to equip households, offices, or shops for purposes such as storage, seating, sleeping faldstool, airline seat, bicycle parking rack 388 933
body of water Any significant accumulation of water, generally on a planet’s surface dammed lake, deep-sea hydrothermal vent, marshland 379 792
weather State of the atmosphere cold snap, tropical cyclone, sea of fog 151 304
precipitation Liquid or solid water that falls to the ground hail, thunderstorm, snowfall 43 72

### v3

@@ F Querying entities with SPARQL
B-16 $\star$ Only organisms 11.3 10.2 51.4 39.2 46.1 62.1 25.9 35.1 26.8
B-32 EntityNet 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 No organism 9.0 22.5 39.2 28.0 33.6 37.7 29.1 32.0 7.5
B-32 Only organisms 4.1 10.2 36.0 21.0 31.5 39.6 8.1 17.6 8.1
B-32 No attributes 8.7 21.8 54.8 33.8 47.9 49.2 24.4 36.9 10.5
B-32 50% alt text 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 100% alt text 13.1 32.7 59.1 38.1 51.1 55.6 30.1 40.8 13.1
B-32 0% alt text 13.1 32.7 55.7 35.5 48.0 53.0 27.1 36.8 12.4
B-32 Full size 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 1/2 size 6.6 16.4 54.1 33.6 47.2 51.0 24.4 36.5 8.9
B-32 1/4 size 3.3 8.2 45.2 25.7 39.1 39.9 18.1 25.6 5.9
B-32 1/8 size 1.6 4.1 33.3 17.7 28.5 28.7 12.0 15.3 4.2
B-32 1/16 size 0.8 2.0 19.9 10.0 17.2 17.6 6.8 5.7 2.8
B-32 Batch size 2k 13.1 32.7 59.7 40.4 52.5 59.4 31.3 44.5 14.4
B-32 Batch size 4k 13.1 32.7 60.9 41.0 53.0 60.0 32.1 44.9 14.9
B-32 Batch size 8k 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 Batch size 16k 13.2 32.7 60.5 39.7 53.2 58.0 30.6 42.9 13.8
B-32 Batch size 32k 13.3 32.7 58.6 37.2 51.3 54.6 27.9 40.6 11.6
Table 10: Hyperparameters used for pretraining and finetuning, unless otherwise stated. For all experiments we use the AdamW [53] optimizer with $\epsilon=1e-8$, $\beta_{1}=0.9$, $\beta 2=0.98$.
Dataset Model Batch Learning Weight Epochs Warmup
Size Rate Decay epochs
CC12M ViT-B/32 8k 5e-4 0.2 18 2
CC12M ViT-B/16 8k 5e-4 0.2 18 2
Ours, pretraining ViT-B/32 8k 5e-4 0.2 18 2
Ours, pretraining ViT-B/16 8k 5e-4 0.2 18 2
Ours, finetuning ViT-B/32 32k 5e-5 0.2 18 2
Ours, finetuning ViT-B/16 32k 5e-5 0.2 18 2
@@
[…]
fontfamily=cmtt,formatcom=,,
PREFIX wdt: <http://www.wikidata.org/prop/direct/>PREFIX wd: <http://www.wikidata.org/entity/>PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>PREFIX schema: <http://schema.org/>PREFIX wikibase: <http://wikiba.se/ontology#>PREFIX skos: <http://www.w3.org/2004/02/skos/core#>SELECT DISTINCT ?ent ?label ?desc ?links (GROUP_CONCAT(DISTINCT ?alias; SEPARATOR=";;;") AS ?aliases)WHERE { VALUES ?typ { wd:Q42889 } ?ent wdt:P279* ?typ . ?ent rdfs:label ?label . FILTER(LANG(?label) = "en") ?ent ˆschema:about/wikibase:sitelinks ?links . FILTER(?links >= 5) OPTIONAL { ?ent schema:description ?desc . FILTER(LANG(?desc) = "en") } OPTIONAL { ?ent skos:altLabel ?alias . FILTER(LANG(?alias) = "en") }}GROUP BY ?ent ?label ?desc ?linksORDER BY DESC(?links)
@@
Table 12: Vehicle entities and accompanying information as extracted from the Wikidata knowledge graph. Showing the first 5 and last 5 out of 17,015 entities. Note that we only collect entities with sitelinks $\geq 5$. The corresponding SPARQL query is shown in Figure˜5.
@@
Q1420 motor car motorized road vehicle designed to carry one to eight people rather than primarily goods 237 auto, motor vehicle, motor cars, motorcar, cars, car, automobiles, automobile, autocar
Q11442 bicycle pedal-driven two-wheel vehicle 203 bike, Bicycles, cycle, pushbike, pedal cycle, pedal bike
Q197 airplane powered fixed-wing aircraft 196 airplane, aeroplane, plane, powered fixed-wing aircraft, planes, plane, aeroplane, fixed-wing powered aircraft, fixed-wing airplane, aeroplanes, fixed-wing aeroplane, airplanes
Q870 train form of rail transport consisting of a series of connected vehicles 193 rail-train, trains, railway train, railtrain, rail train, railroad train
Q11446 ship large buoyant watercraft 178 marine vessel, vessel, water vessel, ships
@@
Q7077241 Odakyu 20000 series RSE Japanese electric multiple unit trainset 5 RSE, Romancecar RSE, Resort Super Express, Odakyu Romancecar RSE, 20000 series
Q812263 Bavarian Pt 2/3 class of 97 German 2-4-0T locomotives 5 ÖBB 770, DR Class 70.0, DRG Class 70.0
@@
Q812260 Bavarian PtL 2/2 class of 6+29+13 German 0-4-0T locomotives 5 DB Class 98.3, DRG Class 98.3, ÖBB 688
@@
product Anything that can be offered to a market banh mi, navigation system 63,676 144,715
substance Any composed matter whose origin is either biological, chemical, or mineral solid lubricant, Chinese tea 34,259 111,383
physical tool Physical item that can be used to achieve a goal Patient lift, police transport 32,727 71,227
animal Kingdom of multicellular eukaryotic organisms saw-scaled viper, Sporathraupis cyanocephala 28,000 76,408
plant Living thing in the kingdom of photosynthetic eukaryotes Whitebark Pine, Eucalyptus coccifera 28,000 55,925
material Substance that can occur in different amounts, all with some similar [mixture of some] characteristics, and with which objects can be made dietary proteins, stone slab tomb 18,021 40,822
vehicle Mobile machine used for transport, whether it has an engine or not, including wheeled and tracked vehicles, air-, water-, and space-craft shipwrecks (objects), Evergreen A-class container ship 17,015 37,849
geographical feature Components of planets that can be geographically located hydrothermal Vents, grooves 8,683 19,030
food Any substance consumed to provide nutritional support for the body; form of energy stored in chemical coffee milk, tikka 8,464 15,332
architectural structure Human-designed and -made structure rock temples, summerhouse 4,507 10,354
anatomical structure Entity with a single connected inherent 3d shape that’s created by coordinated expression of the organism’s own dna bronchi, maxillary wisdom tooth 4,394 9,999
facility Place, equipment, or service to support a specific function public toilet, automobile servicing shop 2,767 6,740
physical activity Human physical activity consisting of voluntary bodily movement by skeletal muscles American rules football, archery 2,228 4,422
clothing Covering worn on the body blucher shoe, G-suit 1,929 4,313
building Structure, typically with a roof and walls, standing more or less permanently in one place shoestore, family restaurant 1,655 3,964
musical instrument Device created or adapted to make musical sounds electroencephalophone, Chinese flutes 1,450 3,493
organ Collection of tissues with similar functions nasal bone, cranial nerves 1,155 2,450
furniture Movable objects used to equip households, offices, or shops for purposes such as storage, seating, sleeping faldstool, airline seat 388 933
body of water Any significant accumulation of water, generally on a planet’s surface dammed lake, deep-sea hydrothermal vent 379 792
weather State of the atmosphere cold snap, tropical cyclone 151 304
precipitation Liquid or solid water that falls to the ground hail, thunderstorm 43 72